OpenStudy (anonymous):

12/k=24/k k=1.4?????

OpenStudy (anonymous):

?

OpenStudy (anonymous):

solve the proportion

OpenStudy (anonymous):

Well, the only value that makes that 'true' is k=0.

OpenStudy (anonymous):

\[\frac{12}{k}=\frac{24}{k} \rightarrow \frac{k}{12}=\frac{k}{24} \rightarrow 2k=k \rightarrow k=0\]

OpenStudy (anonymous):

12k=24k 12k-24k=0 -12k=0 k=0/12 k=0

OpenStudy (anonymous):

\[\frac{12}{k}=\frac{24}{k} \] Move the fraction on the right hand side of the equation to the left hand side. Combine the fractions and simplify. The result is: \[-\frac{12}{k}==0 \] If the fraction is zero then 12 has to be equal to zero. By the way, WolframAlfa.com says "there is no solution" which goes without say I guess.